# ELU-GCN: Effectively Label-Utilizing Graph Convolutional Network

### v1

@@ 2.3 Graph Contrastive Learning
In the ELU graph $\mathbf{S}^{*}$, all nodes are theoretically ELU nodes. However, the original graph $\widehat{\mathbf{A}}$ includes ELU nodes and NELU nodes. Obviously, in representation learning, the representations of ELU nodes in both $\mathbf{S}^{*}$ and $\widehat{\mathbf{A}}$ should be consistent for keeping common information related to the class, the representations of NELU nodes in $\widehat{\mathbf{A}}$ should be different from their representation in $\mathbf{S}^{*}$. To do this, we first propose to learn a projection head $p_{\theta}$ to map both the ELU graph representations and the original graph representations into the same latent space, i.e., $\overline{\mathbf{P}}=p_{\theta}(\overline{\mathbf{H}})$ and $\widetilde{\mathbf{P}}=p_{\theta}(\widetilde{\mathbf{H}})$, where $\overline{\mathbf{H}}$ is the representation of the output layer of the GCN dominated by the original graph, and $\widetilde{\mathbf{H}}$ is the representation of the output layer of the GCN dominated by the ELU graph. We then design a contrastive loss as follows:
$\mathcal{L}_{\mathrm{con}}=$ $-\log\frac{\frac{1}{|V_{\mathrm{ELU}}|}\sum^{V_{\mathrm{ELU}}}_{i=0}\exp(d(\overline{\mathbf{P}}_{i},\widetilde{\mathbf{P}}_{i})/\tau)}{\frac{1}{|V_{\mathrm{ELU}}|}\sum^{V_{\mathrm{ELU}}}_{i=0}\exp(d(\overline{\mathbf{P}}_{i},\widetilde{\mathbf{P}}_{i})/\tau)+\frac{1}{|V_{\mathrm{NELU}}|}\sum^{V_{\mathrm{NELU}}}_{j=0}\exp(d(\overline{\mathbf{P}}_{j},\widetilde{\mathbf{P}}_{j})/\tau)}$ (10)
$+\gamma\log\left(\sum_{i,j=1}^{d}e^{\overline{\mathbf{P}}^{T}\overline{\mathbf{P}}+\widetilde{\mathbf{P}}^{T}\widetilde{\mathbf{P}}}\right)$
where $d(\cdot)$ is distance function, $\tau$ denotes the temperature parameter and $\gamma$ is a hyper-parameter.
In Eq. (10), the first term encourages minimizing the distance between every ELU node in the ELU graph and its corresponding node in the original graph, while maximizing the distance between every NELU node in the original graph and its corresponding node in the ELU graph. The second term ensures that different dimensions of the representation matrices (i.e., $\overline{\mathbf{P}}$ and $\widetilde{\mathbf{P}}$) are uniformly distributed over the latent space, thereby avoiding the issue of feature collapse. As a result, Eq. (10) is available to extract the consistency and mutually exclusive information between the representations dominated by the ELU graph and the original graph.
@@
$\mathcal{L}=CE((1-\eta)Softmax(\overline{\mathbf{H}})+(\eta)Softmax(\widetilde{\mathbf{H}}),\mathbf{Y})+\lambda\mathcal{L}_{con}$ (11)
where $\eta,\lambda\in[0,1]$ are hyper-parameters to fuse the predicted results of two views and two objective functions, respectively.
@@ 2.4 Theoretical Analysis
The ELU graph has been shown to effectively utilize the label information in Section 2.1. In this section, we theoretically analyze that the generalization ability of the GCN is related to the graph structure and the training labels by Theorem 2.3 (The proof can be found in Appendix B.4):
@@ 3 Experiments
In this section, we conduct experiments on eleven public datasets to evaluate the proposed method (including citation networks, Amazon networks, social networks, and web page networks), compared to structure improvement methods are shown in Appendix F. Additional experimental results are shown in Appendix G.
@@ 3.1 Experimental Setup
Table 1: Performance on node classification task. The highest results are highlighted in bold. "OOM" denotes out of memory.
@@ 3.1.3 Evaluation Protocol
To evaluate the effectiveness of the proposed method, we follow the commonly used setting. Specifically, for the citation network (i.e., Cora, Citeseer, and Pubmed), we use the public split recommended by (Kipf & Welling, 2017) with fixed 20 nodes per class for training, 500 nodes for validation, and 1000 nodes for testing. For Social networks (i.e., Caltech, UF, Hamilton, and Tulane), we randomly generate different data splits with an average train/val/test split ratio of 60%/20%/20%. For the Webpage network (i.e., Chameleon, Squirrel) and co-purchase networks (i.e., Computers, Photo), we all use the public splits recommended in the original papers.
3.2 Rsults Analysis
3.2.1 Effectiveness Analysis
@@
We first evaluate the effectiveness of the proposed method by reporting the results of node classification in Table 1 and Appendix G, respectively. Obviously, the proposed method obtains better performance on seven datasets than comparison methods.
@@
First, compared with traditional GNN methods and advanced GNN methods. the proposed ELU-GCN outperforms them by large margins on most datasets. For example, the proposed ELU-GCN on average improves by 4.05 %, compared to GCN, and improves by 3.26 % compared to the best advanced GCN method (i.e., PCNet), on all datasets. This demonstrates the superiority of graph structure learning methods, as the label information cannot be effectively utilized for many nodes in the original graph. Second, compared to the improvement methods, the proposed ELU-GCN achieves the best results, followed by GSR, GCN-LPA, CoGSL, PTDNet-GCN, NeuralSparse-GCN, and NodeFormer. For example, our method on average improves by 2.21% compared to the best comparison method GSR on all seven datasets. This can be attributed to the fact that the proposed ELU-GCN, which can obtain a graph structure (i.e., the ELU graph) that is more suitable for the GCN model to effectively utilize the label information and efficiently mine the consistency and mutually exclusive information between the original graph and the newly obtained graph. In addition, the Webpage networks (i.e., Chameleon and Squirrel) are heterophilic graphs. As mentioned in the theoretical analysis section, the original graph is difficult to guarantee the generalization ability of GCN, especially for heterophilic graphs. Experimental results show that the proposed ELU-GCN outperforms the GCN using the original heterophilic graph by an average of 9.5%, confirming the results of our theoretical analysis. Consequently, the effectiveness of the proposed method is verified in node classification tasks.
@@
We further evaluate the effectiveness of the proposed method on social network datasets and report the results of node classification in Appendix G.1. We can observe that the proposed method also achieves competitive results on the social network datasets compared to other baselines. For example, the proposed method outperforms the best baseline (i.e., GSR), on almost all datasets.
Table 2: Ablation study.
@@
+$\mathcal{L}_{con}$ 84.04$\pm$0.39 73.17$\pm$0.62 80.51$\pm$0.21 83.72$\pm$2.31 90.80$\pm$1.33 70.59$\pm$1.76 60.91$\pm$1.81
3.2.2 Ablation Study
@@
According to Table 2, we can draw the following conclusions. First, our proposed method achieves the best performance when each component is present, indicating that each is essential. This demonstrates the importance of both learning the ELU graph and extracting information from the original graph, as they not only enable GCN to effectively utilize labels but also retain important information in the original graph. Second, the ELU graph component provided the biggest improvement. For example, the ELU graph improves performance by an average of 2.9% compared to not considering it, and the $\mathcal{L}_{con}$ term improves performance by an average of 1.3% compared to not considering it. This illustrates the importance of enabling nodes to effectively utilize the label information.
3.2.3 Visualization
To verify the effectiveness of the learned ELU graph, we visualize the adjacency matrix of the ELU graph in the heatmap on the Cora, Computers, Photo, and Chameleon datasets and report the results in Figure 3.
Specifically, the rows and columns of heatmaps are reordered by node labels. In the heatmaps, the lighter a pixel, the larger the value of the ELU graph matrix weight. From Figure 3, we observe that the heatmaps exhibit a clear block diagonal structure, with each block corresponding to a category. This indicates that the obtained ELU graph tends to increase the weight connections between nodes of the same category and avoid noisy connections from different classes. As a result, the training labels will be transferred to nodes of the same category under the GCN framework with a high probability, thereby reducing intra-class variance and increasing inter-class distance. Especially on the Chameleon dataset, where the original graph tends to connect nodes with different labels with a high probability (i.e., heterophily). Fortunately, our method can still obtain a graph structure where nodes are connected with the same category, as shown by the experimental results, demonstrating the universality of our method.
@@ 3.2.3 Visualization
Figure 3: Visualization of the adjacency matrix of the ELU graph on Cora, Computers, Photo, and Chameleon datasets.
3.2.4 Generalization Gap Analysis
(a) G.G. of GCN on Cora
(b) G.G. of ours on Cora
(c) G.G. of GCN on Citeseer
(d) G.G. of ours on Citeseer
Figure 4: Visualization of the generalization gap (i.e., G.G) of our model (i.e., ELU-GCN) and GCN on Cora and Citeseer datasets.
As Theorem 2.4 mentioned the proposed ELU-GCN can ensure the generalization ability but GCN using the original graph cannot (i.e., traditional GCN). Therefore, to verify the generalization ability, we introduce the generalization gap (Keskar et al., 2016), which is the gap between the training loss and validating loss. A small gap between the two losses indicates a model with good generalization. We visualize the generalization gap of ELU-GCN and GCN on Cora and Citeseer datasets, the results are shown in Figure 4.
Specifically, the proposed ELU-GCN shows a small generalization gap, compared to GCN. For example, the proposed method’s generalization gap on the Cora and Citeseer datasets is approximately 63.6% and 26.7% lower than that of GCN, respectively. This is consistent with the observation in Theorem 2.4 and further verifies the effectiveness of the proposed ELU-GCN.
4 Running Time V.S. Accuracy
The biggest limitation of graph structure learning methods is the need to query in $\mathcal{O}(n^{2})$ space when learning graph structures. Although we have previously analyzed that the complexity of the proposed algorithm in graph construction is $\mathcal{O}(nc^{3})$ ($c^{3}\ll n$), plus the final graph structure is $\mathcal{O}(n^{2})$ (only one calculation is required), we further test the overall actual running time of the proposed ELU-GCN and compared with the commonly used baseline (i.e., GCN and GAT). The results are in Figure 8.
From Figure 8, we have the observations as follows. First, the overall running time of the proposed method is slightly inferior to GCN, but significantly ahead of GAT. Second, the proposed method achieves the best classification performance. Combining the above two points, the proposed method achieves the optimal trade-off between running time and model performance.
Figure 5: Running time V.S. accuracy.
@@ 5 Conclusion
In this paper, we study the label utilization of GCN and reveal that a considerable number of unlabeled nodes cannot effectively utilize label information in the GCN framework. Furthermore, we propose a standard for determining which unlabeled nodes can effectively utilize label information in the GCN framework. To make more nodes to effectively utilize label information. We propose an effective label-utilizing graph convolutional network framework. To do this, we optimize the graph structure by constraining every node effectively using label information. Moreover, we design a novel contrastive loss to minimize consistency or mutually exclusive information between the original graph and the ELU graph. Our theoretical analysis demonstrates that ELU-GCN provides superior generalization capabilities compared to conventional GCNs. Extensive experimental results further validate that our method consistently outperforms state-of-the-art methods.
@@ Appendix A Complexity
$\mathbf{Q}^{(i)}$ $=\mathbf{H}(\mathbf{Q}^{(i-1)})^{T}\left(\frac{1}{\beta}\mathbf{I}_{N}-\frac{1}{\beta^{2}}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}\right)\mathbf{Q}^{(i-1)}$ (12)
$=\mathbf{H}(\mathbf{Q}^{(i-1)})^{T}\left(\frac{1}{\beta}\mathbf{Y}-\frac{1}{\beta^{2}}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}\mathbf{Q}^{(i-1)}\right).$
We first let $\mathbf{B}=\frac{1}{\beta^{2}}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}\mathbf{Q}^{(i-1)}$ and compute it from right to left. Specifically, the matrix inversion operation on a $c\times c$ matrix is $\mathcal{O}(c^{3})$. Therefore, the overall time complexity of $\mathbf{S}\in\mathbb{R}^{n\times c}$ is $\mathcal{O}(nc^{2}+c^{3})$, where $c\ll n$. Then we can compute $\mathbf{H}(\mathbf{Q}^{(i-1)})^{T}\mathbf{B}$, likewise, we calculate it from right to left, this can reduce the time complexity from $\mathcal{O}(n^{2}c)$ to $\mathcal{O}(nc^{2})$. Therefore the overall time complexity of calculating Eq. LABEL:eq:10 is $\mathcal{O}(nc^{2}+c^{3})$. This significantly improves the model efficiency.
@@
$\mathbf{S}^{*}=\mathbf{H}(\mathbf{Q}^{(i)})^{T}\left(\mathbf{H}\mathbf{H}^{T}+\beta\mathbf{I}_{N}\right)^{-1}=\mathbf{H}(\mathbf{Q}^{(i)})^{T}\left(\frac{1}{\beta}\mathbf{I}-\frac{1}{\beta^{2}}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}\right).$ (13)
@@
$\mathbf{S}^{*}$ $=\mathbf{H}(\mathbf{Q}^{(i)})^{T}\left(\frac{1}{\beta}\mathbf{I}-\frac{1}{\beta^{2}}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}\right)$ (14)
$=\mathbf{H}(\frac{1}{\beta}(\mathbf{Q}^{(i)})^{T}-\frac{1}{\beta^{2}}(\mathbf{Q}^{(i)})^{T}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T})$
We first let $\mathbf{P}=\frac{1}{\beta^{2}}(\mathbf{Q}^{(i)})^{T}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}$ and calculate $(\mathbf{Q}^{(i)})^{T}\mathbf{H}$, wich time complexity is $\mathcal{O}(nc^{2})$, then we can get a $c\times c$ matrix $(\mathbf{Q}^{(i)})^{T}\mathbf{H}$, the time complexity of $\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}$ is $\mathcal{O}(c^{3})$, thus the overall complexity of $\mathbf{P}$ is $\mathcal{O}(nc^{2}+c^{3})$. Finally, the complexity of $\mathbf{HP}$ is $\mathcal{O}(n^{2}c)$, since $c$ is the number of classes, it have $c\ll n$. Therefore, the complexity grows quadratically with the number of samples i.e., $\mathcal{O}(n^{2})$.

### v3

@@ 3.3 Contrastive Constraint
In the ELU graph $\mathbf{S}^{*}$, all nodes are theoretically ELU nodes. However, the original graph $\widehat{\mathbf{A}}$ includes ELU nodes and NELU nodes. Obviously, in representation learning, the representations of ELU nodes in both $\mathbf{S}^{*}$ and $\widehat{\mathbf{A}}$ should be consistent for keeping common information related to the class, the representations of NELU nodes in $\widehat{\mathbf{A}}$ should be different from their representation in $\mathbf{S}^{*}$.
To do this, we first propose to learn a projection head $p_{\theta}$ to map both the ELU graph representations and the original graph representations into the same latent space, i.e., $\overline{\mathbf{P}}=p_{\theta}(\overline{\mathbf{H}})$ and $\widetilde{\mathbf{P}}=p_{\theta}(\widetilde{\mathbf{H}})$, where $\overline{\mathbf{H}}$ is the representation of the output layer of the GCN dominated by the original graph, and $\widetilde{\mathbf{H}}$ is the representation of the output layer of the GCN dominated by the ELU graph, the output of the model is $\widehat{\mathbf{Y}}=\mathrm{Softmax}((1-\eta)\overline{\mathbf{H}}+\eta\widetilde{\mathbf{H}})$, where $\eta$ is a hyper-parameter. We then design a contrastive loss as follows:
$\mathcal{L}_{\mathrm{con}}$ $=-\log\frac{\mathrm{Pos}}{\mathrm{Pos}+\mathrm{Neg}}-\frac{1}{n}\sum^{n}_{i=1}\sum^{c}_{j=1}\hat{y}_{i,j}\log\hat{y}_{i,j}$ (12)
$s.t.\left\{\begin{matrix}\mathrm{Pos}=\frac{1}{|V_{\mathrm{ELU}}|}\sum^{V_{\mathrm{ELU}}}_{i=0}\exp(d(\overline{\mathbf{P}}_{i},\widetilde{\mathbf{P}}_{i})/\tau)\\ \mathrm{Neg}=\frac{1}{|V_{\mathrm{NELU}}|}\sum^{V_{\mathrm{NELU}}}_{j=0}\exp(d(\overline{\mathbf{P}}_{j},\widetilde{\mathbf{P}}_{j})/\tau)\end{matrix}\right.$
where $d(\cdot)$ is the distance function, $\tau$ denotes the temperature parameter, and $\hat{y}_{i,j}$ is the element in row $i$ and column $j$ in $\widehat{\mathbf{Y}}$.
In Eq. (12), the first term encourages minimizing the distance between every ELU node in the ELU graph and its corresponding node in the original graph, while maximizing the distance between every NELU node in the original graph and its corresponding node in the ELU graph. The second term encourages the decision boundary to be positioned in low-density regions, enhancing the distinction between nodes and ensuring the underlying assumption of semi-supervised learning (Berthelot et al., 2019). As a result, Eq. (12) is available to extract the consistency and mutually exclusive information between the representations dominated by the ELU graph and the original graph.
@@
$\mathcal{L}=CE(\widehat{\mathbf{Y}},\mathbf{Y})+\lambda\mathcal{L}_{con}$ (13)
where $\lambda\in[0,1]$ is a hyperparameter to fuse the predicted results of two views and two objective functions.
@@ 3.4 Theoretical Analysis
The ELU graph has been shown to enable the GCN to effectively utilize label information, as demonstrated in Section 3.1. In this section, we theoretically analyze that the generalization ability of the GCN is related to the graph structure and the training labels by Theorem 3.3 (The proof can be found in Appendix B.4):
@@ 4 Experiments
In this section, we conduct experiments on eleven public datasets to evaluate the proposed method (including citation networks, Amazon networks, social networks, and web page networks), compared to structure improvement methods111 The code is released at https://github.com/huangJC0429/label-utilize-GCN. Detailed settings are shown in Appendix E. Additional experimental results are shown in Appendix F.
@@ 4.1 Experimental Setup
Table 1: Performance on node classification task. The highest results are highlighted in bold. ”OOM” denotes out of memory.
@@ 4.1.3 Evaluation Protocol
To evaluate the effectiveness of the proposed method, we follow the commonly used setting. Specifically, for the citation network (i.e., Cora, Citeseer, and Pubmed), we use the public split recommended by (Kipf & Welling, 2017) with fixed 20 nodes per class for training, 500 nodes for validation, and 1000 nodes for testing. For Social networks (i.e., Caltech, UF, Hamilton, and Tulane), we randomly generate different data splits with an average train/val/test split ratio of 60%/20%/20%. For the Webpage network (i.e., Chameleon, Squirrel) and co-purchase networks (i.e., Computers, Photo), we use the public splits recommended in the original papers.
4.2 Effectiveness Analysis
@@
We first evaluate the effectiveness of the proposed method by reporting the results of node classification in Table 1 and Appendix F, respectively. Obviously, the proposed method obtains better performance on seven datasets than comparison methods.
First, compared with traditional GNN methods and advanced GNN methods. the proposed ELU-GCN outperforms them by large margins on most datasets. For example, the proposed ELU-GCN on average improves by 4.05 %, compared to GCN, and improves by 3.26 % compared to the best advanced GCN method (i.e., PCNet), on all datasets. This demonstrates the superiority of graph structure learning methods, as the label information cannot be effectively utilized for many nodes in the original graph.
@@
Second, compared to the improvement methods, the proposed ELU-GCN achieves the best results, followed by GSR, GCN-LPA, CoGSL, PTDNet-GCN, NeuralSparse-GCN, and NodeFormer. For example, our method on average improves by 2.21% compared to the best comparison method GSR on all seven datasets. This can be attributed to the fact that the proposed ELU-GCN, which can obtain a graph structure (i.e., the ELU graph) that is more suitable for the GCN model to effectively utilize the label information and efficiently mine the consistency and mutually exclusive information between the original graph and the newly obtained graph. In addition, the Webpage networks (i.e., Chameleon and Squirrel) are heterophilic graphs. As mentioned in the theoretical analysis section, the original graph is difficult to guarantee the generalization ability of GCN, especially for heterophilic graphs. Experimental results show that the proposed ELU-GCN outperforms the GCN using the original heterophilic graph by an average of 9.5%, confirming the results of our theoretical analysis. Consequently, the effectiveness of the proposed method is verified in node classification tasks.
@@
We further evaluate the effectiveness of the proposed method on social network datasets and report the results of node classification in Appendix F.1. We can observe that the proposed method also achieves competitive results on the social network datasets compared to other baselines. For example, the proposed method outperforms the best baseline (i.e., GSR), on almost all datasets.
Table 2: Classification performance of each component in the proposed method on all datasets.
@@
+$\mathcal{L}_{con}$ 84.29$\pm$0.39 74.23$\pm$0.62 80.51$\pm$0.21 83.73$\pm$2.31 90.81$\pm$1.33 70.90$\pm$1.76 56.91$\pm$1.81
4.3 Ablation Study
@@
According to Table 2, we can draw the following conclusions. First, our proposed method achieves the best performance when each component is present, indicating that each is essential. This demonstrates the importance of both learning the ELU graph and extracting information from the original graph, as they not only enable GCN to effectively utilize labels but also retain important information in the original graph.
Second, the ELU graph component provided the biggest improvement. For example, the ELU graph improves performance by an average of 2.9% compared to not considering it, and the $\mathcal{L}_{con}$ term improves performance by an average of 1.3% compared to not considering it. This illustrates the importance of unlabeled nodes being affected by effectively labeled information in message passing.
@@
Figure 3: Visualization of the adjacency matrix of the ELU graph on Cora, Computers, Photo, and Chameleon datasets. The rows and columns are nodes that are reordered based on node labels, the lighter a pixel, the larger the value of the ELU graph matrix weight.
4.4 Visualization
To provide an intuitive and clear understanding of the effectiveness of the learned ELU graph, we visualize the adjacency matrix of the ELU graph in the heatmap on the Cora, Computers, Photo, and Chameleon datasets and report the results in Figure 3.
Specifically, the rows and columns of heatmaps are reordered by node labels. In the heatmaps, the lighter a pixel, the larger the value of the ELU graph matrix weight. From Figure 3, we observe that the heatmaps exhibit a clear block diagonal structure, with each block corresponding to a category. This indicates that the obtained ELU graph tends to increase the weight of connections between nodes of the same category and avoid noisy connections from different classes. As a result, under the GCN framework, the training nodes are likely to propagate the label information to unlabeled nodes of the same category with a high probability, thereby reducing intra-class variance and increasing inter-class distance. Especially on the Chameleon dataset, where the original graph tends to connect nodes with different labels with a high probability (i.e., heterophily). Fortunately, our method can still obtain a graph structure where nodes are connected with the same category, as shown by the experimental results, demonstrating the universality of the proposed method.
4.5 Analyze the Trade-off Between Running Time and Accuracy
The biggest limitation of graph structure learning methods is the need to query in $\mathcal{O}(n^{2})$ space when learning the adjacency matrix of graph structures. In our proposed method, we cleverly leverage the inverse matrix transformation trick to avoid the computational complexity of $\mathcal{O}(n^{2})$ or even higher. Although we have previously analyzed that the complexity of the proposed algorithm in graph construction is $\mathcal{O}(nc^{3})$ ($c^{3}\ll n$), plus the final graph structure is $\mathcal{O}(n^{2})$ (only one calculation is required), we further test the overall actual running time and accuracy of the proposed ELU-GCN and compared with the commonly used baseline (i.e., GCN and GAT). The results are in Figure 4.
Figure 4: Scatter plot showing the relationship between model runtime and accuracy, where the x-axis represents the runtime of different models on different datasets and the y-axis represents their corresponding accuracy (%).
From Figure 4, we have the observations as follows. First, the overall running time of the proposed method is slightly inferior to GCN, but significantly ahead of GAT. This indicates that the proposed ELU graph learning method does not incur too much time overhead and is comparable to the basic GNN model. Second, the proposed method achieves the best classification performance. Combining the above two points, the proposed method achieves the optimal trade-off between running time and model performance.
@@ 5 Conclusion
In this paper, we study the label utilization of GCN and reveal that a considerable number of unlabeled nodes cannot effectively utilize label information in the GCN framework. Furthermore, we propose a standard for determining which unlabeled nodes can effectively utilize label information in the GCN framework. To enable more unlabeled nodes to utilize label information effectively. We propose an effective label-utilizing graph convolutional network framework. To do this, we optimize the graph structure following the above standard, enabling every unlabeled node to effectively leverage label information. Moreover, we design a novel contrastive loss to capture consistency or mutually exclusive information between the original graph and the ELU graph. Our theoretical analysis demonstrates that ELU-GCN provides superior generalization capabilities compared to conventional GCNs. Extensive experimental results further validate that our method consistently outperforms state-of-the-art methods.
Acknowledgement
Xiaofeng Zhu was supported in part by the National Key Research and Development Program of China under Grant (No.2022YFA1004100). Xiaoshuang Shi was supported in part by the Sichuan Science and Technology Program under Grant (No. 2024YFHZ0268).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Complexity
$\mathbf{Q}^{(i)}$ $=\mathbf{Q}^{(i-1)}\mathbf{H}^{T}\left(\frac{1}{\beta}\mathbf{I}_{N}-\frac{1}{\beta^{2}}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}\right)\mathbf{Q}^{(i-1)}$ (14)
$=\mathbf{Q}^{(i-1)}\mathbf{H}^{T}\left(\frac{1}{\beta}\mathbf{Y}-\frac{1}{\beta^{2}}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}\mathbf{Q}^{(i-1)}\right).$
We first let $\mathbf{B}=\frac{1}{\beta^{2}}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}\mathbf{Q}^{(i-1)}$ and compute it from right to left. Specifically, the matrix inversion operation on a $c\times c$ matrix is $\mathcal{O}(c^{3})$. Therefore, the overall time complexity of $\mathbf{S}\in\mathbb{R}^{n\times c}$ is $\mathcal{O}(nc^{2}+c^{3})$, where $c\ll n$. Then we can compute $\mathbf{Q}^{(i-1)}\mathbf{H}^{T}\mathbf{B}$, likewise, we calculate it from right to left, this can reduce the time complexity from $\mathcal{O}(n^{2}c)$ to $\mathcal{O}(nc^{2})$. Therefore the overall time complexity of calculating Eq. LABEL:eq:10 is $\mathcal{O}(nc^{2}+c^{3})$. This significantly improves the model efficiency.
@@
$\mathbf{S}^{*}=\mathbf{Q}^{(i-1)}\mathbf{H}^{T}\left(\mathbf{H}\mathbf{H}^{T}+\beta\mathbf{I}_{N}\right)^{-1}=\mathbf{Q}^{(i-1)}\mathbf{H}^{T}\left(\frac{1}{\beta}\mathbf{I}-\frac{1}{\beta^{2}}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}\right).$ (15)
@@
$\mathbf{S}^{*}$ $=\mathbf{Q}^{(i-1)}\mathbf{H}^{T}\left(\frac{1}{\beta}\mathbf{I}-\frac{1}{\beta^{2}}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}\right)$ (16)
$=\mathbf{Q}^{(i-1)}(\frac{1}{\beta}\mathbf{H}^{T}-\frac{1}{\beta^{2}}\mathbf{H}^{T}\mathbf{H}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T})$
We first let $\mathbf{P}=\frac{1}{\beta^{2}}\mathbf{H}\mathbf{H}^{T}\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}\mathbf{H}^{T}$ and calculate $\mathbf{H}^{T}\mathbf{H}$, wich time complexity is $\mathcal{O}(nc^{2})$, then we can get a $c\times c$ matrix $\mathbf{H}^{T}\mathbf{H}$, the time complexity of $\left(\mathbf{I}_{c}+\frac{1}{\beta}\mathbf{H}^{T}\mathbf{H}\right)^{-1}$ is $\mathcal{O}(c^{3})$, thus the overall complexity of $\mathbf{P}$ is $\mathcal{O}(nc^{2}+c^{3})$. Finally, the complexity of $\mathbf{Q}^{(i-1)}\mathbf{P}$ is $\mathcal{O}(n^{2}c)$, since $c$ is the number of classes, it have $c\ll n$. Therefore, the complexity grows quadratically with the number of samples i.e., $\mathcal{O}(n^{2})$.